# Physics-informed, boundary-constrained Gaussian process regression for the reconstruction of fluid flow fields

### v2

@@ 1 Introduction
In order to completely define a GP, a mean function and a covariance function (or a kernel as defined in the dedicated literature) are required. These functions define the mathematical properties of the prior estimate, thus yielding an automatic satisfaction of the problem constraints if desired. The nature of these constraints can be set by physical specifications of the modelling problem, still no unique rule-of-thumb exists for incorporating them in the mean and/or kernel functions. Specific techniques to learn them from the data have been proposed in e.g. [1, 36]. Divergence-free random vector fields have been considered in [43] and used for Lagrangian flow particle simulation in [34]. Additionally, interests in constraining GP priors to boundary conditions on the boundary of domains with different geometries have also emerged [12, 19, 45]. In [19], the boundary-constrained GPs are built from a particular structure based on Fourier expansions, whereas [45] presents a spectral framework based on eigenfunctions of the Laplacian that is applicable to stationary base kernels. Other methods, such as [12], define boundary-constrained GP kernels from multiplication of the base kernel with a symmetric function satisfying the boundary conditions. Our chosen method, developed firstly by [17], modifies any given GP with a continuous kernel to satisfy a null condition (also known as homogeneous Dirichlet condition in the context of partial differential equations (PDEs)) over any compact subset of its domain. We provide an associated numerical method for scalability of this procedure under different choices of the probability measure. To our knowledge, this is the first time that such a general boundary-constraining method (adapted for computing kernel derivatives) is applied to an arbitrary continuous kernel (not only stationary) and under a complex definition of the arbitrary constraining set, such as the boundary of a NACA airfoil.
@@ 2 Reminder on Gaussian Process Regression
Here ${G}({\boldsymbol{X}},{\boldsymbol{X}})$ is the Gram matrix with elements ${G}({\boldsymbol{X}},{\boldsymbol{X}})]_{ij}={G}({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{j})$ and ${G}({\boldsymbol{x}},{\boldsymbol{X}})$ is the column vector with elements ${G}({\boldsymbol{x}},{\boldsymbol{X}})_{i}={G}({\boldsymbol{x}},{\boldsymbol{x}}_{i})$. Moreover, the counterpart for quantifying the covariance of the random variables ${Z}({\boldsymbol{x}})$ and ${Z}({\boldsymbol{x}}^{\prime})$ associated to the locations $\boldsymbol{x}$} and $\boldsymbol{x}$$^{\prime}$ in the domain $\Omega$}, conditioned on ${Z}({\boldsymbol{X}})={\boldsymbol{Y}}$, reads:
@@ 3.4 Benefits and limitations with respect to discrete constraints
From a more general perspective, when performing regression by means of formulas (2.1) and (2.3), the most expensive computational task is the evaluation and inversion of the corresponding Gram matrices. For instance, given a set of $N$ observations ${\boldsymbol{Y}}=\smash{(y_{1},\ldots,y_{N})^{\mathsf{T}}}\in\mathbb{R}^{N}$ corresponding to points ${\boldsymbol{X}}=\smash{\left\{{\boldsymbol{x}}_{i}\right\}_{i=1}^{N}}$ inside $\Omega$}, consider the representation formula:
@@
Inversion of Gram matrix (using Cholesky decomposition) $\approx\displaystyle\smash{\frac{1}{3}}N^{3}$ $\approx\displaystyle\smash{\frac{1}{3}}(N+{N_{\Gamma}})^{3}$
Table 1: Comparison of the complexity of two approaches for the enforcement of the Dirichlet boundary condition (3.6) on the posterior distribution from the base GP Z{Z}. As a reference, we consider the complexity of a single evaluation of kernel G{G} of order O(1)(1).
@@ 4 Physics-informed GPR for incompressible flows in aerodynamics
This section outlines the approach for constructing physics-informed GPR in the context of fluid dynamics. We particularly address incompressible fluid flows past an aerodynamic profile as described in Sect. 4.1. In Sect. 4.2 we focus on how the divergence-free condition pertaining to incompressible fluid flows can be imposed to GP prior definitions, while simultaneously constraining them with the condition imposed along the profile boundary by means of BCGPs. The overall GPR strategy we have developed is fully detailed in Sect. 4.3. Particular aspects for computing BCGP kernel derivatives as needed in this setting are discussed in Sect. 4.4, while some considerations on the use of non-constrained kernels are diverted to Sect. 4.5 with the aim to compare this more classical approach to the proposed one.
@@ 4.1 Fluid dynamics context
4.2 Physics-informed, boundary-constrained GP
Now we present our approach of modelling an incompressible fluid flow in the setting of the previous section. The main ingredient is to model the velocity stream function $\psi$} as a scalar GP with a prior definition that includes the slip boundary condition on $\Gamma$ and the divergence-free condition on $\Omega$}.
More precisely, we consider an arbitrary GP Z{Z} indexed on the spatial domain $\Omega$} with a covariance kernel G{G} belonging to $\mathcal{C}^{3}({\Omega}\times{\Omega})$. Since the profile boundary $\Gamma$ is a compact set inside $\Omega$}, we model $\psi$} by a BCGP defined from this base GP Z{Z}. That is, we consider ${\psi}={{Z}_{0}}\sim{\mathcal{GP}}(0,{{G}_{0}})$ as the BCGP associated to Z{Z} through the procedure of Proposition (3.1), for ${G}_{0}$} being the constrained kernel defined in Eq. (3.10). It follows that the flow velocity can be modelled as the zero-mean vector-valued GP:
@@ 4.2 Physics-informed, boundary-constrained GP
${\textbf{curl}}\,{{Z}_{0}}({\boldsymbol{x}})\cdot{\boldsymbol{n}}({\boldsymbol{x}})=O,\quad\text{for all}\,{\boldsymbol{x}}\in\Gamma\,$, (4.7)
@@ 4.3 GPR observation design
where ${\boldsymbol{X}}={{\boldsymbol{X}}_{\text{b}}}\cup{{\boldsymbol{X}}_{\text{d}}}$ is the expanded set of total observation positions and ${\boldsymbol{V}}(t)=\smash{({\boldsymbol{V}}_{\text{b}}(t)^{\mathsf{T}},{\boldsymbol{V}}_{\text{d}}(t)^{\mathsf{T}})}^{\mathsf{T}}$ is the $\mathbb{R}^{2N}$ column vector of the corresponding velocity information, with $N=N_{\text{b}}+{N_{{\Omega}}}$. This reconstruction formula allows us to approximate the fluid flow velocity at any unobserved point $\boldsymbol{x}$} in the domain $\Omega$} while inherently satisfying the incompressibility and the profile boundary conditions continuously at all points, respectively. A representation formula for ${{\boldsymbol{u}}^{\star}}(\cdot,t)$ can be established along the same lines as in Sect. 2, so that:
${{\boldsymbol{u}}^{\star}}({\boldsymbol{x}},t)={\textbf{curl}}_{\boldsymbol{x}}\otimes{\textbf{curl}}_{{\boldsymbol{x}}^{\prime}}\,{{G}_{0}}({\boldsymbol{x}},{\boldsymbol{X}})\,{\boldsymbol{\beta}_{0}}(t)$ (4.9)
@@
${\boldsymbol{\beta}_{0}}(t)=\left({\textbf{curl}}_{\boldsymbol{x}}\otimes{\textbf{curl}}_{{\boldsymbol{x}}^{\prime}}\,{{G}_{0}}({\boldsymbol{X}},{\boldsymbol{X}})+\eta{\boldsymbol{I}}\right)^{-1}{\boldsymbol{V}}(t)\,.$ (4.10)
This representation considers a regularization parameter $\eta>0$ (the so-called nugget). A covariance counterpart for computing the correlation between estimates from two points inside $\Omega$} can be established similarly to Eq. (2.2) or Eq. (2.4).
@@
whereas the vorticity field, defined as:
@@ 4.4 Derivatives of a constrained kernel
4.5 Discrete boundary constraints
Alternatively to the reconstruction formulas in the previous Sect. 4.3 using a BCGP, a different approach for enforcing the slip boundary condition of Eq. (4.2) can be established in a discrete setting as follows. Let us consider design values ${\boldsymbol{V}}_{\text{b}}$ and ${\boldsymbol{V}}_{\text{d}}$ for the velocity at positions ${\boldsymbol{X}}_{\text{b}}$} and ${\boldsymbol{X}}_{\text{d}}$}, corresponding to outer domain boundary and inner domain observations, respectively. In addition, consider a discretization of the profile boundary ${{\boldsymbol{X}}_{\Gamma}}=\smash{\{{\boldsymbol{x}}^{i}_{\Gamma}\}_{i=1}^{N_{\Gamma}}}$. Then, by modelling the velocity stream function $\psi$} directly from the base GP ${Z}\sim{\mathcal{GP}}(0,{G})$, we can consider the following approach for the reconstruction of the velocity field:
where $\boldsymbol{x}$} is any unobserved point in $\Omega$}. In this approach, the boundary constraint of Eq. (4.2) is enforced in a discrete way in the posterior GP distribution. A representation formula can be derived in the same way as for Eq. (4.9), and similar reconstruction formulas hold. This approach will be used in the next section as a means of comparison with the BCGP-based approach.
@@ 5.1 Profiles
    $\gamma_{1}(s)=x(s)-\operatorname{sgn}(\sin s)\frac{{y_{\text{t}}}(x(s))y_{\text{c}}^{\prime}(x(s))}{\sqrt{\left[y_{\text{c}}^{\prime}(x(s))\right]^{2}+1}}\,,\quad\gamma_{2}(s)=y_{\text{c}}(x(s))+\operatorname{sgn}(\sin s)\frac{{y_{\text{t}}}(x(s))}{\sqrt{\left[y_{\text{c}}^{\prime}(x(s))\right]^{2}+1}}\,.$ (5.1)
@@ 5.3 Flow around a cylinder profile
A snapshot of the ground-truth velocity field is shown in Figure 3 for a fixed time instant. The GPR reconstruction results are presented only for this time instant.
@@
5.3.1 Continuous profile boundary constraints
@@
Next, we present the GPR reconstruction of the velocity field using our physics-informed framework by adding observation points inside the computational domain. That is, we consider the reconstruction formula (4.9) with ${\boldsymbol{X}}={{\boldsymbol{X}}_{\text{b}}}\cup{{\boldsymbol{X}}_{\text{d}}}$ and ${\boldsymbol{V}}(t)=\smash{({\boldsymbol{V}}_{\text{b}}(t)^{\mathsf{T}},{\boldsymbol{V}}_{\text{d}}(t)^{\mathsf{T}})}^{\mathsf{T}}$ as in Sect. 4.3. A total of $N=415$ observation points of the velocity was considered in the computational domain $\Omega$}, of which $N_{\text{b}}=88$ and ${N_{{\Omega}}}=327$ observations belong to $\partial{\Omega}$ and $\Omega$}, respectively. As before, no design observation on the profile boundary $\Gamma_{\text{cyl}}$} is included. In order to compute the spectral expansion for the definition of $\smash{{{G}_{0}^{(I,J)}}}$ from Eq. (3.23) and (4.19), a uniform regular discretization of ${\mathcal{P}}=[0,2\pi]$ in $I=400$ intervals and the largest $J+1=19$ eigenvalues of the expansion were used (see below). The interpolation of the velocity field is presented in Figure 6.
@@
5.3.2 Discrete boundary constraints
We now present a comparison of the BCGP-based approach with the discrete setting of Sect. 4.5, while keeping a constant budget of design observations for both cases.
To do this, we fix this budget in terms of the size of the Gram matrix used for interpolation, say ${N_{\text{Gram}}}=830$. Then, we give specific configurations of the design observations to each approach. For the BCGP-based method, we consider the same configuration of $N=415$ observations of the velocity as in Figure 6. In the case of the discrete approach, we consider $N=344$ observations of the velocity inside the domain $\Omega$} and its outer boundary, and we add ${N_{\Gamma}}=142$ observation points for the consideration of the slip boundary condition in Eq. (4.20). The velocity field estimates from physics-informed GPR with ${G}_{0}$} and with G{G} using discrete boundary constraints are presented in Figure 8. The relative error fields corresponding to each estimate, with respect to the ground truth from Figure 3, are also depicted. Furthermore, the configuration of the GPR design for each approach is summed up in Table 2, including the relative normal fit indicator $\epsilon_{n}$ of Eq. (4.15). It is observed that the BCGP-based method allows to satisfy the slip boundary condition at a better order of accuracy than that of the discrete method for the same observation budget. Moreover, this budget allows to include observations of the velocity at other parts of the domain instead of adding discrete points on the profile boundary.

### v3

@@ 1 Introduction
In order to completely define a GP, a mean function and a covariance function (or a kernel as defined in the dedicated literature) are required. These functions define the mathematical properties of the prior estimate, thus yielding an automatic satisfaction of the problem constraints if desired. The nature of these constraints can be set by physical specifications of the modelling problem, still no unique rule-of-thumb exists for incorporating them in the mean and/or kernel functions. Specific techniques to learn them from the data have been proposed in e.g. [1, 38, 20]. Divergence-free random vector fields have been considered in [45] and used for Lagrangian flow particle simulation in [36]. Additionally, interests in constraining GP priors to boundary conditions on the boundary of domains with different geometries have also emerged [12, 19, 47]. In [19], the boundary-constrained GPs are built from a particular structure based on Fourier expansions, whereas [47] presents a spectral framework based on eigenfunctions of the Laplacian that is applicable to stationary base kernels. Other methods, such as [12], define boundary-constrained GP kernels from multiplication of the base kernel with a symmetric function satisfying the boundary conditions. Our chosen method, developed firstly by [17], modifies any given GP with a continuous kernel to satisfy a null condition (also known as homogeneous Dirichlet condition in the context of partial differential equations (PDEs)) over any compact subset of its domain. We provide an associated numerical method for scalability of this procedure under different choices of the probability measure. To our knowledge, this is the first time that such a general boundary-constraining method (adapted for computing kernel derivatives) is applied to an arbitrary continuous kernel (not only stationary) and under a complex definition of the arbitrary constraining set, such as the boundary of a NACA airfoil.
@@ 2 Reminder on Gaussian Process Regression
Here ${G}({\boldsymbol{X}},{\boldsymbol{X}})$ is the Gram matrix with elements ${G}({\boldsymbol{X}},{\boldsymbol{X}})]_{ij}={G}({\boldsymbol{x}}_{i},{\boldsymbol{x}}_{j})$ and ${G}({\boldsymbol{x}},{\boldsymbol{X}})$ is the row vector with elements ${G}({\boldsymbol{x}},{\boldsymbol{X}})_{i}={G}({\boldsymbol{x}},{\boldsymbol{x}}_{i})$. Moreover, the counterpart for quantifying the covariance of the random variables ${Z}({\boldsymbol{x}})$ and ${Z}({\boldsymbol{x}}^{\prime})$ associated to the locations $\boldsymbol{x}$} and $\boldsymbol{x}$$^{\prime}$ in the domain $\Omega$}, conditioned on ${Z}({\boldsymbol{X}})={\boldsymbol{Y}}$, reads:
@@ 3.4 Benefits and limitations with respect to discrete constraints
From a more general perspective, when performing regression by means of formulas (2.1) and (2.3), the most expensive computational task is the evaluation and inversion of the corresponding Gram matrices. For instance, given a set of $N$ observations ${\boldsymbol{Y}}=\smash{(y_{1},\ldots,y_{N})^{\mathsf{T}}}\in\mathbb{R}^{N}$ corresponding to the locations ${\boldsymbol{X}}=\smash{\left\{{\boldsymbol{x}}_{i}\right\}_{i=1}^{N}}$ inside $\Omega$}, consider the representation formula:
@@
Inversion of Gram matrix $\approx\displaystyle\smash{\frac{1}{3}}N^{3}$ $\approx\displaystyle\smash{\frac{1}{3}}(N+{N_{\Gamma}})^{3}$
Table 1: Comparison of the complexity of two approaches for the enforcement of the Dirichlet boundary condition (3.6) on the posterior distribution from the base GP Z{Z}. As a reference, we consider the complexity of a single evaluation of kernel G{G} of order O(1)(1). The inversion of the Gram matrix is performed by Cholesky decomposition.
@@ 4 Physics-informed GPR for incompressible flows in aerodynamics
This section outlines the approach for constructing physics-informed GPR in the context of fluid dynamics. We particularly address incompressible fluid flows past an aerodynamic profile as described in Sect. 4.1. In Sect. 4.2 we focus on how the divergence-free condition pertaining to incompressible fluid flows can be imposed to GP prior definitions, while simultaneously constraining them with the condition imposed along the profile boundary by means of BCGPs. The overall GPR strategy we have developed is fully detailed in Sect. 4.4. Particular aspects for computing BCGP kernel derivatives as needed in this setting are discussed in Sect. 4.5, while some considerations on the use of non-constrained kernels are diverted to Sect. 5.3.2 with the aim to compare this more classical approach to the proposed one.
@@ 4.1 Fluid dynamics context
4.2 Divergence-free, boundary-constrained GPs
We now outline the GP framework to model flow fields while accounting for the divergence-free and boundary conditions continuously. A natural choice is to model the velocity stream function $\psi$} as a scalar GP with a prior definition that includes the slip boundary condition on $\Gamma$ and the divergence-free condition within $\Omega$}.
More precisely, we consider an arbitrary GP Z{Z} indexed on the spatial domain $\Omega$} with a covariance kernel G{G} belonging to $\mathcal{C}^{3}({\Omega}\times{\Omega})$. Since the profile boundary $\Gamma$ is a compact set inside $\Omega$}, we model $\psi$} by a BCGP defined from this base GP Z{Z}. That is, we consider ${\psi}={{Z}_{0}}\sim{\mathcal{GP}}(0,{{G}_{0}})$ as the BCGP associated to Z{Z} through the procedure of Proposition 3.1, for ${G}_{0}$} being the constrained kernel defined in Eq. (3.10). It follows that the flow velocity can be modelled as the zero-mean vector-valued GP:
@@ 4.2 Divergence-free, boundary-constrained GPs
${\textbf{curl}}\,{{Z}_{0}}({\boldsymbol{x}})\cdot{\boldsymbol{n}}({\boldsymbol{x}})=O,\quad\forall{\boldsymbol{x}}\in\Gamma\,$, (4.7)
@@
4.3 Energy decay GPs
The next step is to define the structure of the base kernel G{G}, from which the boundary-constrained kernel ${G}_{0}$} is computed. Since ${G}\in\mathcal{C}^{3}({\Omega}\times{\Omega})$ must hold, one choice is to define it as an anisotropic radial-basis function (RBF):
${G}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\sigma^{2}\exp\left[-\frac{(x_{1}-x^{\prime}_{1})^{2}+\alpha(x_{2}-x^{\prime}_{2})^{2}}{2\ell^{2}}\right],\quad{\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}\,$, (4.8)
where $\sigma^{2},\ell$, and $\alpha$ are hyperparameters for variance, correlation length and anisotropy, respectively. However, the dynamics of the flow can evolve at different spatial scales. Therefore the kernel of the GP prior should include a multi-scale additive structure to follow the energy decay of these dynamics, as proposed in [36]. In particular, the expectation of the squared velocity increment from a point ${\boldsymbol{x}}\in{\Omega}$ to ${\boldsymbol{x}}+{\boldsymbol{h}}\in{\Omega}$ is approximately of order $\smash{\left\|{\boldsymbol{h}}\right\|^{2}}$. Thus the underlying structure shall recover this power-law.
Instead of using solely (4.8), we rather define G{G} as an additive kernel with different orders of correlation lengths:
${G}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\sum_{m=0}^{M}\sigma_{m}^{2}\exp\left[-\frac{(x_{1}-x^{\prime}_{1})^{2}+\alpha_{m}^{2}(x_{2}-x^{\prime}_{2})^{2}}{2\ell_{m}^{2}}\right],\quad{\boldsymbol{x}},{\boldsymbol{x}}^{\prime}\in{\Omega}\,$, (4.9)
where $M+1$ is the number of scales. We set the correlation lengths to $\smash{\ell_{m}}=\smash{\ell_{0}/2^{3m}}$ for $m=0,1,\ldots,M$, in order to capture the small spatial scales. The anisotropy parameters are set to $\smash{\alpha_{m}}=\smash{\max\left(1\,,\,2^{3-m}\right)}$ since the horizontal and vertical dynamics are different in magnitude at the coarser scales. This definition satisfies the condition $\alpha_{m}\to 1$ in order to recover an isotropic power-law at the smaller scales. The variance of each scale is set using a decay parameter $\gamma\in\mathbb{N}$, such that $\smash{\sigma_{m}}=\smash{\sigma_{0}/2^{\gamma m}}$. The kernel (4.9) is stationary and thus, by the same arguments as in [36, Sect. 5.2], we have that:
$\mathbb{E}\left\{\ \left\|{\textbf{curl}}\,{Z}({\boldsymbol{x}}+{\boldsymbol{h}})-{\textbf{curl}}\,{Z}({\boldsymbol{x}})\right\|^{2}\ \right\}=\text{O}(\left\|{\boldsymbol{h}}\right\|^{\gamma/3})\,,\quad\left\|{\boldsymbol{h}}\right\|\to 0\,.$
We thus set $\gamma=6$ in order to fit the power-law of the squared velocity increment in two-dimensions. In this way, the prior (and posterior) behavior of the GP is further informed by the energy decay. A similar version of this decay structure is found when modeling the vorticity field directly as a scalar GP, as detailed in [21].
@@ 4.4 GPR observation design
where ${\boldsymbol{X}}={{\boldsymbol{X}}_{\text{b}}}\cup{{\boldsymbol{X}}_{\text{d}}}$ is the expanded set of total observation positions and ${\boldsymbol{V}}(t)=\smash{({\boldsymbol{V}}_{\text{b}}(t)^{\mathsf{T}},{\boldsymbol{V}}_{\text{d}}(t)^{\mathsf{T}})}^{\mathsf{T}}$ is the $\mathbb{R}^{2N}$ column vector of the corresponding velocity information, with $N=N_{\text{b}}+{N_{{\Omega}}}$. This reconstruction formula allows us to approximate the fluid flow velocity at any unobserved point $\boldsymbol{x}$} in the domain $\Omega$} while inherently satisfying the divergence-free and profile boundary conditions continuously at all locations, respectively. A representation formula for $\smash{{{\boldsymbol{u}}^{\star}}(\cdot,t)}$ can be established along the same lines as in Sect. 2 so that:
${{\boldsymbol{u}}^{\star}}({\boldsymbol{x}},t)={{\boldsymbol{K}}_{0}}({\boldsymbol{x}},{\boldsymbol{X}})\,{\boldsymbol{\beta}_{0}}(t)\,$, (4.11)
@@
${\boldsymbol{\beta}_{0}}(t)=\left({{\boldsymbol{K}}_{0}}({\boldsymbol{X}},{\boldsymbol{X}})+\eta{\boldsymbol{I}}\right)^{-1}{\boldsymbol{V}}(t)\,$, (4.12)
and covariance matrix $\smash{{{\boldsymbol{K}}_{0}}}$ given by (4.5). This representation introduces the regularization parameter $\eta>0$ (the so-called nugget) such that the formula (4.11) is actually no longer strictly interpolatory. The covariance counterpart for computing the posterior correlation between estimates from two locations ${\boldsymbol{x}},{\boldsymbol{x}}^{\prime}$ inside $\Omega$} can also be established as:
${\boldsymbol{\Sigma}^{\star}}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})=\operatorname{Cov}\left\{{{\textbf{curl}}\,{{Z}_{0}}}({\boldsymbol{x}}),\,{{\textbf{curl}}\,{{Z}_{0}}}({\boldsymbol{x}}^{\prime})\ |\ {{\textbf{curl}}\,{{Z}_{0}}}({\boldsymbol{X}})={\boldsymbol{V}}(t)\right\}\,.$
By using a matrix-valued formula similar to Eq. (2.2), this last expression becomes:
${\boldsymbol{\Sigma}^{\star}}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})={{\boldsymbol{K}}_{0}}({\boldsymbol{x}},{\boldsymbol{x}}^{\prime})-{{\boldsymbol{K}}_{0}}({\boldsymbol{x}},{\boldsymbol{X}})\left({{\boldsymbol{K}}_{0}}({\boldsymbol{X}},{\boldsymbol{X}})+\eta{\boldsymbol{I}}\right)^{-1}{{\boldsymbol{K}}_{0}}({\boldsymbol{X}},{\boldsymbol{x}}^{\prime})\,.$ (4.13)
@@
whereas the vorticity field $\omega$}, defined by:
@@ 5.1 Profiles
    $\gamma_{1}(s)=x(s)-\operatorname{sgn}(\sin s)\frac{{y_{\text{t}}}(x(s))y_{\text{c}}^{\prime}(x(s))}{\sqrt{\left[y_{\text{c}}^{\prime}(x(s))\right]^{2}+1}}\,,\quad\gamma_{2}(s)=y_{\text{c}}(x(s))+\operatorname{sgn}(\sin s)\frac{{y_{\text{t}}}(x(s))}{\sqrt{\left[y_{\text{c}}^{\prime}(x(s))\right]^{2}+1}}\,$, (5.1)
@@ 5.3 Flow around a cylinder profile
A snapshot of the ground-truth velocity field is shown in Figure 3. The GPR reconstruction results are presented at this iteration.
@@
5.3.1 Continuous enforcement of the boundary conditions on the profile
@@
Next, we present the GPR reconstruction of the velocity field using our physics-informed framework by adding observation points inside the computational domain. That is, we consider the reconstruction formula (4.11) with ${\boldsymbol{X}}={{\boldsymbol{X}}_{\text{b}}}\cup{{\boldsymbol{X}}_{\text{d}}}$ and ${\boldsymbol{V}}(t)=\smash{({\boldsymbol{V}}_{\text{b}}(t)^{\mathsf{T}},{\boldsymbol{V}}_{\text{d}}(t)^{\mathsf{T}})}^{\mathsf{T}}$ as in Sect. 4.4. A total of $N=415$ observation points of the velocity was considered in the computational domain $\Omega$}, of which $N_{\text{b}}=88$ and ${N_{{\Omega}}}=327$ observations belong to $\partial{\Omega}$ and $\Omega$}, respectively. As before, no design observation on the profile boundary $\Gamma_{\text{cyl}}$} is included. In order to compute the spectral expansion for the definition of $\smash{{{G}_{0}^{(I,J)}}}$ from Eq. (3.23) and (4.22), a uniform regular discretization of ${\mathcal{P}}=[0,2\pi]$ in $I=400$ intervals and the largest $J+1=19$ eigenvalues of the expansion were used (see below). The interpolation of the velocity field is presented in Figure 6.
@@
5.3.2 Discrete enforcement of the boundary conditions on the profile
Alternatively to the reconstruction formulas in Sect. 4.4 using a BCGP, a different approach for enforcing the slip boundary condition of Eq. (4.2) can be considered in a discrete setting as follows. Let ${\boldsymbol{V}}_{\text{b}}$ and ${\boldsymbol{V}}_{\text{d}}$ be the observed velocity values at positions ${\boldsymbol{X}}_{\text{b}}$} and ${\boldsymbol{X}}_{\text{d}}$}, corresponding to the outer domain boundary and inner domain observations, respectively. In addition, consider a discretization of the profile boundary ${{\boldsymbol{X}}_{\Gamma}}=\smash{\{{\boldsymbol{x}}^{i}_{\Gamma}\}_{i=1}^{N_{\Gamma}}}$. Then, by modelling the velocity stream function $\psi$} directly from the base GP ${Z}\sim{\mathcal{GP}}(0,{G})$, we can consider the following approach for the reconstruction of the velocity field:
where $\boldsymbol{x}$} is any unobserved point in $\Omega$}. In this approach, the boundary constraint of Eq. (4.2) is enforced in a discrete way in the posterior GP distribution.
We now compare the BCGP-based approach and the discrete approach, while keeping a constant budget of observations for both cases. To do this, we fix this budget in terms of the size of the Gram matrix used for interpolation, say ${N_{\text{Gram}}}=830$. Then, we give specific configurations of the design observations to each approach. For the BCGP-based method, we consider the same configuration of $N=415$ observations of the velocity as in Figure 6. In the case of the discrete approach, we consider $N=344$ observations of the velocity inside the domain $\Omega$} and its outer boundary, and we add ${N_{\Gamma}}=142$ observation points for the consideration of the slip boundary condition in Eq. (5.5). The velocity field estimates from physics-informed GPR with ${G}_{0}$} and with G{G} using discrete boundary constraints are presented in Figure 8. The relative error fields corresponding to each estimate, with respect to the ground truth from Figure 3, are also depicted. Furthermore, the configuration of the GPR design for each approach is summed up in Table 2, including the relative normal fit indicator $\epsilon_{n}$ of Eq. (4.18). It is observed that the BCGP-based method allows to satisfy the slip boundary condition at a better order of accuracy than that of the discrete method for the same observation budget. Moreover, this budget allows to include observations of the velocity at other parts of the domain instead of adding discrete points on the profile boundary.